OpenStudy (anonymous):

Rationalize the numerator. a)\[\sqrt{2+h}+\sqrt{2-h}/h\] b)\[xsqrt{x}-8/x-4\]

ganeshie8 (ganeshie8):

the key thing is to use the below identiy cleverly : \[(a+b)(a-b) = a^2 - b^2\]

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

do you know what the "conjugate" is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply by the conjugate

OpenStudy (anonymous):

i tried, it's not working...

OpenStudy (anonymous):

i.e. multiplly \[\frac{\sqrt{2+h}+\sqrt{2-h}}{h}\times \frac{\sqrt{2+h}-\sqrt{2-h}}{\sqrt{2+h}-\sqrt{2-h}}\]

OpenStudy (anonymous):

what do you get in the numerator?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i need help... :( @ganeshie8

OpenStudy (joannablackwelder):

I don't get 1 for the numerator:\[(2+h) - (2-h)\]

OpenStudy (joannablackwelder):

Is \[2+h-2+h\]

OpenStudy (joannablackwelder):

Which is 2h.

OpenStudy (joannablackwelder):

Does that make sense?

OpenStudy (anonymous):

gotcha, i made stupid mistakes. Thanks for pointing that out! :)

OpenStudy (joannablackwelder):

No worries :)

OpenStudy (anonymous):

do you know how to do b)?

OpenStudy (joannablackwelder):

Is b \[\frac{ x \sqrt{x-8} }{ x-4 }\]

OpenStudy (joannablackwelder):

?

OpenStudy (anonymous):

no, it's \[\frac{ x \sqrt{x}-8 }{ x-4 }\] Sorry of the confusion...

OpenStudy (joannablackwelder):

Ah, gotcha. No worries.

OpenStudy (joannablackwelder):

Let me work it out, just a sec.

OpenStudy (anonymous):

okay thanks!

OpenStudy (joannablackwelder):

This is another multiply by the conjugate problem.

OpenStudy (anonymous):

is this correct?? \[\frac{ x^2+4x+16 }{ x \sqrt{x}+8 }\]

OpenStudy (joannablackwelder):

Hm, that's not what I get.

OpenStudy (joannablackwelder):

Did you do it like this? \[\frac{ x \sqrt{x}-8 }{ x-4 }\times \frac{ x \sqrt{x}+8 }{ x \sqrt{x}+8 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i got \[\frac{ x^3-64 }{ x-4(x \sqrt{x}-8 }\]

OpenStudy (joannablackwelder):

:) That's almost what I get. I just have a + before the 8 in the denominator.

OpenStudy (anonymous):

i meant +8)*

OpenStudy (joannablackwelder):

Ok, cool! :D

OpenStudy (anonymous):

we can reduce that, right?

OpenStudy (joannablackwelder):

Yep.

OpenStudy (joannablackwelder):

Do you know how?

OpenStudy (anonymous):

well, i got \[\frac{ (x-4)(x^2+4x+16) }{ (x-4)(x \sqrt{x}+8) }\] than I reduce the (x-4)

OpenStudy (jhannybean):

I also got :\[\frac{x^3-64}{(x-4)(x\sqrt{x}+8)}\]x\(^3\)-64 is a difference of cubes. \(a^3-b^3 = (a-b)(a^2+ab+b^2)\) This correlates to : \((x)^3 -(4)^3\)

OpenStudy (joannablackwelder):

Perfect!

OpenStudy (jhannybean):

Whoah, faster than me!

OpenStudy (anonymous):

thanks for helping guys! :D @JoannaBlackwelder @Jhannybean

OpenStudy (jhannybean):

No problem :) Don't forget to reduce the like terms in both numerator and denominator!

OpenStudy (joannablackwelder):

You're welcome! :)